# Coherence-Assisted Entanglement Activation During Pure Dephasing Evolution

### v1

@@ 5 Conclusions
We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolution. We extended the previous results  Roszak and Cywiński (2015); Roszak (2018), obtained for pure initial states of the system, to the case of mixed states. We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a lower bound on the generated entanglement, measured by the relative entropy.
Our second result describes another lower bound and upper bound for generated entanglement in terms of the difference of coherence in the system before and after the evolution. These bounds also can be seen as an extension of the result obtained in Streltsov et al. (2015). Although both the lower bound and upper bound contains the extra terms other than decoherence, we successfully prove that larger entanglement indicates more decoherence in pure dephasing evolution with arbitrary environmental state and vice versa.
Regarding the first result, its lower bound character is a price to pay for the generalization to mixed states. However, it certifies that at least the given amount of entanglement was created during the evolution. The first result is the strongest for the two-dimensional central system, where we obtained a compact formula. For arbitrary dimension, our bound still contains a certain optimization. As an example, we calculated analytically and studied numerically our bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics. Obviously, more systems can be studied, especially that in some of them the fidelities has already been calculated Korbicz et al. (2014); Tuziemski et al. (2019); Kwiatkowski et al. (2021); Kiciński and Korbicz (2021)

### v2

@@ 5 Conclusions
We studied generation of entanglement between the system of interest and its environment for a broad class of so called pure dephasing evolutions, also known as measurement-type interactions. In this part of the work, we extended previous results  [4, 5] obtained for pure initial states of the system, to the case of mixed states. We combined methods and results from coherence theory, state discrimination, and open systems evolution to obtain a series of bounds on the amount of the generated entanglement as measured by the relative entropy. Firstly, and most importantly, we obtain lower bounds, which certify that at least the given amount of entanglement was created during the evolution out of a given amount of the initial coherence. The lower bound character of this result is a price to pay for the generalization to arbitrary, mixed states of the system. For qubit as the central system, where we obtained a closed expression, connecting amount of entanglement to the state distinguishability of the environmental states. This makes an interesting link to certain studies of quantum Darwinism [9]. For an arbitrary dimension, our bound still contains a certain optimization. As an example, we calculated analytically and studied numerically the lower bound for the popular spin-boson system, rediscovering the well-known traces of the non-Markovianity but in the entanglement dynamics. Obviously, more systems can be analyzed, especially that for some of them the relevant state distinguishability has already been quantified [37, 38, 39, 33].
Next, we derived both lower and upper bounds for the entanglement in terms of the difference of coherence in the system before and after the evolution, which is a measure of the decoherence efficiency. This proves in an elegant and general way that stronger entanglement leads to more decoherence and vice versa, more decoherence is accompanied with more entanglement generation. These bounds also can be seen as an extension of the result obtained in [17].
Summarizing, our results allow to better characterize and understand the correlations that pure dephasing interaction can produce under general initial conditions. Because of the ubiquitous character of the pure dephasing interaction in open systems [2, 3], our results are of a fairly universal character across the field of open systems and decoherence theory. Moreover, since quantum measurement theory is largely based on the same type of interaction, we characterize how much entanglement is generated during a measurement process, thus linking to foundations of quantum mechanics. On the other hand, a possible practical importance of our studies may lie in the development of non-standard decoherence-mitigation techniques, which rely on the system-environment correlations (see e.g. [40] for an example of such a technique).